# ScaleBiO: Scalable Bilevel Optimization for LLM Data Reweighting

### v1

@@ 3.2 Fully First-order Hypergradient Method
Note that the inner-level problem of Problem (1) can be viewed as a constraint of the outer-level problem, and thus it is tempting to reformulate
@@
Here an auxiliary variable $u$ is introduced to transform the inner problem $w_{\ast}(\lambda)=\arg\min_{w}L_{2}(\lambda,w)$ to be the constraint $L_{2}(\lambda,w)-\min_{u}L_{2}(\lambda,u)=0$ where $u$ serves as the proxy of $w_{\ast}(\lambda)$. By penalizing the constraint of (3.2) with multiplier $\alpha>0$ into the outer-level objective, we have the minimax formulation (Kwon et al., 2023; Lu and Mei, 2023):
$\min_{\lambda\in\Lambda,w}\max_{u}\,\,\mathcal{L}^{\alpha}(\lambda,w,u)$ (5)
@@
To precisely describe the optimality of the minimax problem with the stationarity of the bilevel problem, we overload the notation in (5) and define
@@
$w_{\ast}^{\alpha}(\lambda)$ $:=\arg\min_{w}\Phi^{\alpha}(\lambda,w).$ (9)
We make the following assumptions for the proposed minimax problem throughout this paper.
@@
Under Assumption 1, as indicated by Lemma 1, if $\alpha$ goes to infinity, the stationary point of the minimax problem (5) is also a stationary point of the bilevel problem (1).
@@ 3.3 Proposed Algorithm
By adopting the approach in Section 3.2, the outer-inner dependency can be naturally eliminated from the original bilevel problem and reformulated as an equivalent min-max problem. We propose a single-loop framework, referred to as ScaleBiO, to solve the large-scale min-max problem, outlined in Algorithm 1. In many large-scale machine-learning scenarios, the dimensionality of the training variables can be exceptionally high. For bilevel problems, we employ the randomized block coordinate method (Nesterov, 2012; Pan et al., 2024) to update the inner variables $u,w$, which improves memory efficiency. As for the outer variable $\lambda$, we utilize the stochastic gradient descent method for its update. Besides, the popular methods such as SGD momentum and Adam (Kingma and Ba, 2015) can be seamlessly integrated into Algorithm 1 to update each variable $u$, $w$, and $\lambda$. The multiplier $\alpha$ is predefined with a sufficiently large factor that ensures the min-max solution is a good approximation of the original bilevel problem.
@@ 3.4 Theoretical Results
In this part, we provide a convergence analysis of Algorithm 1. Before showing the details of theoretical results, we introduce the notations for partitions. Let $\left\{x^{1},x^{2},\cdots,x^{J}\right\}$ with $x^{j}\in\mathbb{R}^{d_{j}\times 1}$ be $J$ non-overlapping blocks of $x$. Let the matrix $U_{j}\in\mathbb{R}^{d\times d_{j}}$ be $d_{j}$ columns of a $d\times d$ permutation matrix $U$ corresponding to $j$ block coordinates in $x$. For any partition of $x$ and $U$,
@@
We present the essential lemmas in appendix C to show the theoretical properties of minimax objective $\mathcal{L}^{\alpha}$ in (5), as well as its optimizers $u_{\ast}$ and $w_{\ast}^{\alpha}$. Lemma 1 provides clear evidence that $\Gamma^{\alpha}(\lambda)$ is smooth with parameter $\ell_{\Gamma}=\mathcal{O}(\kappa^{3})$ which is independent on the multiplier $\alpha$.
@@ 4 Experiments
Our experiments are two-fold: We conduct (1) Small Scale Experiments in Section 4.1 for the intuitive understanding of our method ScaleBiO and (2) Real-World Application Experiments in Section 4.2 to validate its effectiveness and scalability. We start with detailing our main findings on the basic capabilities of our approach. Then we scale them up in real-world scenarios to address the challenge of finding the optimal sampling weight when fine-tuning LLM with multiple data sources.
@@ 4.1 Small Scale Experiments
To verify the effectiveness of ScaleBiO in small-scale LLMs, we conducted experiments with GPT-2 (124M) on three tasks with synthetic datasets: data denoising, multilingual training, and instruction-following fine-tuning. Full details are available in Appendix B.1.
@@ 4.1.1 Data Denoising
This experiment aims to verify ScaleBiO’s effectiveness in the data-denoising tasks. Our validation dataset, denoted as $\mathcal{D}_{\text{val}}$, comprises 1000 clean samples randomly selected from the Alpaca dataset (Taori et al., 2023). The training dataset, $\mathcal{D}_{\text{trn}}$, is derived from two distinct sources: the first includes 1000 clean samples also from Alpaca, while the second incorporates 9000 samples from Alpaca that have been artificially corrupted with synthetic noise, where the outputs are replaced with ".".
@@ 4.1.2 Multilingual Training
It is also intriguing to check if ScaleBiO can recover optimal sampling weights for more general distributions. To this end, we design multilingual training experiments, where our validation data $\mathcal{D}_{\text{val}}$ comprises 600 random samples from Alpaca-GPT4-ZH (Peng et al., 2023) and 400 random samples from Alpaca-GPT4-EN (Peng et al., 2023). Hence, the underlying optimal weight is 6:4. In contrast, the training set $\mathcal{D}_{\text{trn}}$ has a 1:1 mix ratio, which consists of 40,000/40,000 random examples from Alpaca-GPT4-EN and Alpaca-GPT4-ZH, respectively.
@@ 4.1.3 Instruction Following
In instruction-following fine-tuning tasks, there is a fundamental tradeoff between diversity and quality. To verify if ScaleBiO can deduce these implicit weights of low- and high-quality datasets, we conduct experiments on instruction-following tasks with GPT-2, where Alpaca and Alpaca-GPT4 (Peng et al., 2023) are employed. Here Alpaca-GPT4 shares the same instructions and input as Alpaca, whose high quality is distinguished by its outputs generated from a more sophisticated model GPT-4 (Achiam et al., 2023). Our validation data for bilevel optimization $\mathcal{D}_{\text{val}}$ consists of 1000 random samples from Alpaca-GPT4, while our training data $\mathcal{D}_{\text{trn}}$ consists of 2 separate parts: 1000 random samples from Alpaca-GPT4 and 9000 random samples from Alpaca.
As shown in Figure 3, although Alpaca-GPT4 accounts for only a small proportion of the training data (10%), it is highlighted by our method, revealing that our method effectively up-weights the high-quality data source, leading to improved model outcomes.
@@ 4.2 Real-World Application Experiments
In this section, we take a step forward to land bilevel optimization in real-world data reweighting applications, where we aim to balance multiple training data sources for boosting LLMs’ overall performance in downstream tasks.
4.2.1 Large-Scale Data Reweighting
LLaMA-3-8B LLaMA-3-13B GPT-NeoX-20B Yi-34B
source weight source weight source weight source weight
WildChat 0.711 WildChat 0.711 Airoboros 0.986 ShareGPT4 0.627
Airoboros 0.154 ShareGPT4 0.137 ShareGPT4 0.005 Airoboros 0.111
ChatAlpaca 0.119 ChatAlpaca 0.021 ChatAlpaca 0.003 WildChat 0.105
Total 0.984 Total 0.869 Total 0.994 Total 0.843
Table 2: Data sources with top-3 weights for LLaMA-3-8B, LLaMA-3-13B, GPT-NeoX-20B and Yi-34B.
Figure 4: Learned weights in large-scale data reweighting experiments for Yi-34B. We focus on the top 5 data sources and denote the remaining as Others.
To demonstrate ScaleBiO’s effectiveness in large-scale settings, practical choices of model and datasets are employed under the data reweighting scenarios. Specifically, we collect popular instruction following and chat completion data to construct a diverse training dataset of 18 distinct sources. The training set is formed by 9 high-quality datasets that are mostly from GPT-4 and 9 multilingual datasets serving as less relevant data sources. For our reference dataset $\mathcal{D}_{\text{val}}$, we prompt GPT-4 to generate 1200 two-turn instructions and responses that correspond to the 8 categories in MT-Bench. Our detailed data description and preprocessing can be found in Appendix B.2.
We test our approach on LLaMA-3 model series [^1] https://ai.meta.com/blog/meta-llama-3/ with parameter size of 8B and 13B [^2] https://huggingface.co/Replete-AI/Llama-3-13B. We also conduct experiments on GPT-NeoX-20B (Black et al., 2022) and Yi-34B (01.AI et al., 2024). To evaluate our approach, we first sample 3,000 data from each source for reweighting. Then we sample 10,000 data according to the weights at the end of bilevel optimization to train our model (denotes as Ours). For a fair baseline comparison, we sample 10,000 data uniformly and add the reference data $\mathcal{D}_{\text{val}}$ into them to train the baseline model (denotes as Baseline).
As shown in Table 2 and Figure 4, for models with sizes varied from 8B to 34B, ScaleBiO is capable of filtering out the most irrelevant multi-lingual datasets and putting emphasis on GPT-4 generated data sources. Strikingly, the top three weights in all models were found to be within a small subset of our numerous data sources (WildChat, Airoboros, ChatAlpaca and ShareGPT4), which suggests a universality in the importance attributed to these data sources. Moreover, the top three weights in all models account for over 80% of the total weight. This concentration of importance in a small subset of data sources highlights the effectiveness of our method in identifying and emphasizing the most influential data sources, indicating that our method consistently identifies the most critical data sources, regardless of the model architecture.
4.2.2 Large Scale Instruction Following
To confirm the gain of ScaleBiO, we further apply learned sampling weights to fine-tune LLaMA-3-8B and LLaMA-3-70B models.
We evaluate LLMs’ instruction-following ability using MT-Bench (Zheng et al., 2023a) with single-answer grading. This benchmark prompts chat assistants with challenging multi-turn open-ended questions and uses "LLM-as-a-judge" for evaluation. MT-Bench stands out as a comprehensive benchmark that aligns well with human preference. It comprises 80 questions, evenly distributed across 8 categories: Writing, Roleplay, Extraction, Reasoning, Math, Coding, Knowledge I (STEM), and Knowledge II (humanities/social science). To demonstrate, we compare the performance of LLaMA-3-8B trained with our reweighted data versus the original data. Specifically, we first reweight data sources on LLaMA-3-8B. Then we fune-tune the LLM using the reweighted data.
The results, as shown in Table 3 and Table 4, indicate that our method consistently outperforms the baseline across different categories by a large margin, highlighting the effectiveness of our reweighting approach.
On top of that, the same weights learned from LLaMA-3-8B can be seamlessly transferred to fine-tuning LLaMA-3-70B model. As shown in Table 3, our model achieves noticeable improvement on average compared with the baseline, proving the transferability of the learned data weights across the same family of models.
Model Average
LLaMA-3-8B Baseline 6.11
Ours 7.12
LLaMA-3-8B $\rightarrow$ LLaMA-3-70B Baseline 7.85
Ours 8.05
@@ 4.2.2 Large Scale Instruction Following
4.2.3 Implementation Details
Our data reweighting process lasts for 3 epochs with $\alpha$ equals to 100 and initial learning rate of $10^{-5}$ for models and $10^{-2}$ for weights. For all our fine-tuning processes, we train the LLM for 1 epoch with an initial learning rate of $8\times 10^{-6}$ and a global batch size of 64. Throughout our experiments, we adopt randomized coordinate descent with AdamW (Pan et al., 2024) and bfloat16 precision for efficient training and inference. Our experiments are conducted on 8 NVIDIA A100 80GB GPUs.
@@ 5 Conclusion
In this paper, we propose ScaleBiO, the first bilevel optimization instantiation that is capable of scaling to 34B LLMs on data reweighting tasks. Theoretically, ScaleBiO ensures optimality of the learned data weights and enjoys the same convergence guarantees as conventional first-order penalty-based bilevel optimization algorithms on smooth and strongly convex objectives. Empirically, ScaleBiO enables data reweighting on $\geq 7B$ sized models, bringing forth an efficient data filtering and selection pipeline for improving model performance on various downstream tasks. On top of that, the learned sampling weights on LLaMA-3-8B can be seamlessly transferred to larger-sized LLaMA-3-70B while still obtaining non-trivial performance improvements.

### v2

@@ 3.2 Fully First-order Hypergradient Method
Recent advancements in the theoretical literature of bilevel optimization allow scalable methods to be developed. The main idea is actually quite similar to merging digits in radix sort. The first step is to decouple two “digit” terms and view the inner-level problem in (1) as a higher-order digit,
@@
Here the auxiliary variable $u$ is introduced to detach the inner-outer dependency, which transforms the inner problem $w_{\ast}(\lambda)=\arg\min_{w}L_{2}(\lambda,w)$ to be the constraint $L_{2}(\lambda,w)-\min_{u}L_{2}(\lambda,u)=0$. By prioritizing the “high-order” constraint term of (3.2) with multiplier $\alpha>0$, the minimax formulation in Kwon et al. (2023); Lu and Mei (2023) is recovered:
$\min_{\lambda\in\Lambda,w}\max_{u}\,\,\mathcal{L}^{\alpha}(\lambda,w,u).$ (4)
@@
To precisely describe the optimality of the minimax problem with the stationarity of the bilevel problem, the following notations in (4) are overloaded and defined as
@@
$w_{\ast}^{\alpha}(\lambda)$ $=\arg\min_{w}\Phi^{\alpha}(\lambda,w).$ (8)
Additionally, the following assumptions are made for the proposed minimax problem throughout this paper.
@@
Under Assumption 1, as indicated by Lemma 1, if $\alpha$ goes to infinity, the stationary point of the minimax problem (4) is also a stationary point of the bilevel problem (1). Intuitively, it is like finding a way to the same peak of the mountain with distinct paths, where bilevel optimization suggests a winding road, and minimax utilizes a helicopter.
@@ 3.3 Proposed Algorithm
To solve this reformulated large-scale min-max problem, we introduce ScaleBiO in Algorithm 1, a single-loop framework that is capable of scaling up to 30B-sized models. To further reduce memory consumption, randomized block coordinate methods are employed to update the inner variables $u,w$ (Nesterov, 2012; Pan et al., 2024), where $U_{j},W_{j}\in\mathbb{R}^{d\times d_{j}}$ denotes the block matrices that map the permutation of parameters back to model weights. The optimizer choice varies depending on the backbone model, where Adam or AdamW (Kingma and Ba, 2015; Loshchilov, 2017) is much preferable for LLMs. The penalty coefficient $\alpha$ is predefined with a large factor that ensures the min-max solution is a good approximation of the original bilevel problem.
@@ 3.4 Theoretical Results
In this part, we provide a convergence analysis of Algorithm 1, explaining how fast the algorithm can reach a desired stationary point. Before showing the details of theoretical results, we introduce the notations for partitions. Let $\left\{x^{1},x^{2},\cdots,x^{J}\right\}$ with $x^{j}\in\mathbb{R}^{d_{j}\times 1}$ be $J$ non-overlapping blocks of $x$. Let the matrix $U_{j}\in\mathbb{R}^{d\times d_{j}}$ be $d_{j}$ columns of a $d\times d$ permutation matrix $U$ corresponding to $j$ block coordinates in $x$. For any partition of $x$ and $U$,
@@
The essential lemmas are available in Appendix C to show the theoretical properties of minimax objective $\mathcal{L}^{\alpha}$ in (4), as well as its optimizers $u_{\ast}$ and $w_{\ast}^{\alpha}$. Lemma 1 provides clear evidence that $\Gamma^{\alpha}(\lambda)$ is smooth with parameter $\ell_{\Gamma}=\mathcal{O}(\kappa^{3})$ which is independent on the multiplier $\alpha$.
@@ 4 Experiments
To verify the effectiveness of ScaleBiO, two types of experiment are conducted: (1) Small Scale Experiments in Section 4.1, which offers intuitions for understanding ScaleBiO’s theoretical properties in toy settings, and (2) Real-World Application Experiments in Section 4.2 that validate its scalability in larger-sized models on instruction-following and mathematical reasoning tasks.
@@ 4.1 Small Scale Experiments
To understand the properties of ScaleBiO, experiments with GPT-2 (124M) are conducted on three tasks with synthetic datasets: data denoising, multilingual training, and instruction-following fine-tuning. Full details are available in Appendix B.1.
@@ 4.1.1 Data Denoising
In this experiment, ScaleBiO’s data denoising ability is tested, where noisy samples are expected to be assigned with zero weights. The validation dataset, denoted as $\mathcal{D}_{\text{val}}$, comprises 1000 clean samples randomly selected from the Alpaca dataset Taori et al. (2023). The training dataset, $\mathcal{D}_{\text{trn}}$, is derived from two distinct sources: the first includes 1000 clean samples also from Alpaca, while the second incorporates 9000 samples from Alpaca that have been artificially corrupted with synthetic noise, where the outputs are replaced with ".".
@@ 4.1.2 Multilingual Training
It is also intriguing to check if ScaleBiO can recover optimal sampling weights for more general distributions. To this end, the multilingual training experiments are introduced, where the validation data $\mathcal{D}_{\text{val}}$ comprises 600 random samples from Alpaca-GPT4-ZH Peng et al. (2023) and 400 random samples from Alpaca-GPT4-EN Peng et al. (2023). Hence, the underlying optimal weight is 6:4. In contrast, the training set $\mathcal{D}_{\text{trn}}$ has a 1:1 mix ratio, which consists of 40,000/40,000 random examples from Alpaca-GPT4-EN and Alpaca-GPT4-ZH, respectively.
@@ 4.1.3 Instruction Following
In instruction-following fine-tuning tasks, there is a fundamental tradeoff between diversity and quality. To verify if ScaleBiO can deduce these implicit weights of low- and high-quality datasets, experiments are conducted on instruction-following tasks with GPT-2, where Alpaca and Alpaca-GPT4 Peng et al. (2023) are employed. Here Alpaca-GPT4 shares the same instructions and input as Alpaca, whose high quality is distinguished by its outputs generated from a more sophisticated model GPT-4 (Achiam et al., 2023). The validation data for bilevel optimization $\mathcal{D}_{\text{val}}$ consists of 1000 random samples from Alpaca-GPT4, while the training data $\mathcal{D}_{\text{trn}}$ consists of 2 separate parts: 1000 random samples from Alpaca-GPT4 and 9000 random samples from Alpaca.
As shown in Figure 3, although Alpaca-GPT4 accounts for only a small proportion of the training data (10%), it is highlighted by ScaleBiO, revealing that it can effectively up-weights the high-quality data source, leading to improved model outcomes.
@@ 4.2 Real-World Application Experiments
In this section, ScaleBiO is tested in real-world data reweighting applications, including instruction following and mathematical reasoning tasks, which demonstrates its scalability and empirical benefits in practice.
4.2.1 Instruction Following
Method Model
Llama-3-8B Qwen-2-7B Gemma-2-9B
SOBA (Dagréou et al., 2022) OOM OOM OOM
Uniform Weighting 6.11 6.66 5.31
LESS (Xia et al., 2024b) 6.06 7.18 7.20
RHO-LOSS (Mindermann et al., 2022a) 6.89 7.34 7.38
ScaleBiO 7.12 7.76 7.51
Table 2: Instruction Following. Here all methods are evaluated in MT-Bench (Zheng et al., 2023b) with GPT-4o LLM judge, where scores range from 0 to 10. OOM stands for out of memory.
In the instruction following setting, ScaleBiO is validated under the real-world scenario where the data collection is conducted in a non-filtered fashion, e.g. datasets of weak correlations with the downstream task may be included.
Setup
Three $\sim$7B-sized LLMs, including Llama-3-8B (Dubey et al., 2024), Qwen-2-7B (Yang et al., 2024b), and Gemma-2-9B (Team et al., 2024) are evaluated in the widely adopted benchmark of MT-Bench (Zheng et al., 2023b), where a GPT-4o judge (Hurst et al., 2024) is employed to score the generated responses of each model on 80 high-quality multi-turn questions. Different aspects of the model, such as writing, role play, and STEM, are scored by the GPT-4o judge and averaged in the final MT-Bench score.
The training portfolio comprises $\sim$4.2M total samples from 18 different sources, as detailed in Table 10 in Appendix B.2. All datasets are collected in a task-agnostic fashion, where datasets necessary for general instruction following tasks, but have weak correlations to MT-Bench are also included. One typical example of such datasets is multi-lingual conversations.
To form the training set, all data reweighting methods are required to assign weights to 18 sources and extract 10K samples from the 4.2M portfolio. The target model will be trained on the training set and evaluated to produce the final MT-Bench score.
Results
As shown in Table 2, ScaleBiO is the only bilevel algorithm capable of yielding meaningful weights across data sources. On top of that, SaleBiO outperforms popular influence-aware data filtering method LESS (Xia et al., 2024b) and reference-model-based data reweighting approach RHO-LOSS (Mindermann et al., 2022a), both are considered strong non-bilevel baselines in the data reweighting literature.
4.2.2 Mathematical Reasoning
To further demonstrate ScaleBiO’s data reweighting ability under scenarios with refined dataset sources, a training portfolio similar to Dong et al. (2024) is adopted for mathematical reasoning, where datasets detailed in Table 3 are proven to be conducive to downstream math tasks. Here coding and instruction following datasets are considered necessary, which allow the LLM to learn minimum reasoning and instruction following abilities for answering mathematical questions.
Dataset #Samples
hkust-nlp/dart-math-uniform 591K
Open-Orca/SlimOrca 518K
openbmb/UltraInteract_sft 289K
TIGER-Lab/MathInstruct 262K
microsoft/orca-math-word-problems-200k 200K
WizardLMTeam/WizardLM_evol_instruct_V2_196k 196K
ise-uiuc/Magicoder-Evol-Instruct-110K 110K
anon8231489123/ShareGPT_Vicuna_unfiltered 94K
teknium/GPTeacher-General-Instruct 89K
teknium/GPT4-LLM-Cleaned 55K
Total 2.4M
Table 3: Dataset for Mathematical Reasoning.
Setup
Similar to Section 4.2.1, three models of Llama-3-8B, Qwen-2-7B and Gemma-2-9B are employed. For evaluation, the standard math benchmark of GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b) are utilized. The reweighting methods are expected to extract 20K samples from the given 10 sources to form the training set, with the target model fine-tuned on the set and evaluated to produce the final accuracy.
Method GSM8K (Cobbe et al., 2021) MATH (Hendrycks et al., 2021b)
Llama-3-8B Qwen-2-7B Gemma-2-9B Llama-3-8B Qwen-2-7B Gemma-2-9B
SOBA OOM OOM OOM OOM OOM OOM
Uniform Weighting 53.6 65.0 56.3 14.2 36.7 24.8
RHO-LOSS 53.8 70.7 56.9 13.6 38.8 25.0
LESS 52.5 71.6 57.9 14.0 38.9 28.3
ScaleBiO 56.2 74.1 59.4 15.1 41.7 30.0
Table 4: Mathematical Reasoning. Here all metrics are accuracies ranging from 0 to 100. OOM stands for out of memory.
Results
As shown in Table 4, ScaleBiO consistently outperforms all baselines across different models and benchmarks, by a non-trivial margin of 1%-9%, which demonstrates ScaleBiO’s superiority in reweighting task-oriented datasets.
Method GSM8K MATH
LESS OOM OOM
RHO-LOSS OOM OOM
Uniform Weighting 78.1 54.0
ScaleBiO 87.1 59.8
Table 5: Large-Scale Mathematical Reasoning on Qwen-2.5-32B (Yang et al., 2024a). Here all metrics are accuracies ranging from 0 to 100. OOM stands for out of memory.
To further validate ScaleBiO’s scalability in even larger-sized LLMs, Qwen-2.5-32B (Yang et al., 2024a) is adopted in the same setting. As presented in Table 5, ScaleBiO is the only data reweighting implementation capable of scaling up to this size. Here LESS and RHO-LOSS both run out of GPU memories due to their non-scalable implementation or requirement for extra reference models. In contrast, ScaleBiO has the same space complexity as full parameter fine-tuning, allowing it to be applicable in any single-node training scenarios.
5 Discussion
Existence of Optimal Datasets?
As ScaleBiO is capable of learning optimal task-orient data weights for different models, it serves as a great tool to inspect data weight transferability across different models. As it is unsurprising to find that Llama-3-8B-learned data weights can be transferred to Llama-3-70B and still yield certain improvement (Table 6), it is more intriguing to observe that the learned data weights vary significantly across different model families, as shown in Table 7 of Appendix A.1.
Model MT-Bench score
Llama-3-8B $\rightarrow$ Llama-3-70B Uniform Weighting 7.85
ScaleBiO 8.05
@@ 5 Discussion
It is worth noticing that the weight difference is much smaller inside the same model family. This phenomenon is conjectured to stem from the difference in LLMs’ pre-training dataset distributions, where the strengths of different models may vary from each other and need distinct datasets to adapt to the same downstream task. In that case, optimal dataset weights across models would be impossible for small-sized dataset settings, leaving model-dependent reweighting be the only choice.
@@ 6 Conclusion
In this paper, we propose ScaleBiO, the first bilevel optimization instantiation that is capable of scaling to over 30B-sized LLMs on data reweighting tasks. Theoretically, ScaleBiO ensures optimality of the learned data weights and enjoys the same convergence guarantees as conventional first-order penalty-based bilevel optimization algorithms on smooth and strongly convex objectives. Empirically, ScaleBiO enables data reweighting on $>$30B-sized models, bringing forth an efficient data filtering and selection pipeline for improving model performance on various downstream tasks.